# Bit of a Close Talker: A Practical Guide to Serverless Cloud Co-Location Attacks

### v1

@@ 30 Million requests per year (Coca-Cola’s vending machine service [39]);
    Each function host is allocated 2GB memory and 1GB storage;
In Table II, we provide a comparison of AWS and Azure. For AWS, we provide results from US-East-1 region. We use the AWS Lambda Pricing Calculator [10] to estimate the cost of serverless services, and their EC2 dedicated host pricing information [9] to estimate the cost of having a dedicated host. Similarly, for Azure, we use their dedicated host pricing list [11] and their pricing calculator [12] in the West US region.
@@
Using a dedicated host not only incurs substantially higher costs, but can also lead to scalability limitations. For example, in AWS, a single dedicated host can accommodate only 16 a1-medium instances (based on the capacity information provided in [9]). This level of concurrency is insufficient to handle bursts of requests. In contrast, Double-Dip, as a defense integrated within the serverless framework, preserves the platform’s ability to scale elastically and can therefore support applications with high and variable demand.
@@ X-A Serverless Cloud Scheduling
Scheduler is considered as one of the critical components of a serverless cloud system. In recent years, there have been works that target different features of serverless functions and propose corresponding optimizations to scheduling algorithms. Fuerst et al. [19] proposes an algorithm named CH-RLU, which improves the classic consistent hashing algorithm and takes server loads, function cold-start overheads, etc. into consideration to make trade-offs between server loads and function locality. Kaffes et al. [23] offer a taxonomy of serverless scheduling algorithms and summarize a design parameter space of these schedulers. The authors then propose their algorithm named Hermod, which significantly improves the performance of serverless systems.
Regarding locality optimizations, Aumala et al. [8] propose PASch, which also uses consistent hashing and allows serverless systems to utilize package locality of running functions to accelerate the function start process. Abdi et al. [1] propose to use locality hints in schedulers that specify a function’s preferences in data locality and optimize the performance of serverless functions. Li et al. [28] propose an algorithm to avoid function cold-starts in serverless systems by recycling idle function hosts of other functions and avoiding creating new function hosts.
@@ X-B Co-Location Attacks
The most relevant work is [47]. The authors propose a method to fingerprint physical servers and identify the auto-scaling feature of the Google Cloud Run platform, using which they manage to achieve instance co-location. However, compared to our work, their major focus is on the server fingerprinting step and they only proved that their method works on one type of cloud scheduler. Our work focuses on the systemization of the co-location attack process, and our contribution is the construction of a universal attack strategy that can efficiently target different kinds of serverless cloud schedulers. Also, we propose a low-overhead mitigation scheduling algorithm, which is based on “soft” user isolation, to defend against co-location attacks.
@@ XII Ethics Considerations
The study of co-location itself does not directly compromise cloud users’ instances. It would require subsequent attacks to cause real damages, which are beyond the scope of this work. We disclosed our findings to Microsoft, although they did not classify the issue as a vulnerability, noting that side-channel defenses are already in place. These defenses, however, are orthogonal to our work, which focuses on the scheduler side.
@@ B-A Basics
The placement results of schedulers that consider only resource availability can be approximated as random events in a classical probability model. This can be explained as follows: since cluster information is agnostic to all users, from a user’s perspective, the resource availability on physical servers are random variables following the same distribution. This symmetry indicates that the probability of a user’s instance being placed on any given node is equally $1/N$, where $N$ is the size of the cluster.
@@ B-B Random Scheduling
We first consider the scenario where every invocation to functions is randomly dispatched to nodes in the serverless cluster. In this scenario, each function invocation is considered independently and no locality optimization is taken. If the assigned machine does not have a corresponding warm function host running, it results in a cold-start process. This scheduling policy is load-balanced but suffers low performance due to the lack of locality optimizations.
@@
i.e. co-location tends to happen on almost every physical server if the number of invocations to victim and attacker functions is sufficiently large.
@@ B-C Invocation Locality
Now let’s consider the scenario where the scheduler introduces optimization to avoid cold-start and always assigns function executions to existing function hosts if there are any. In this scenario, we can define $\alpha$ as the number of different attack functions and $\beta$ as the number of launched victim function hosts. In this scenario, we have $\beta\equiv 1$. As there are $\alpha$ different functions, there will be $\alpha$ different function hosts that are placed on physical machines in the cluster.
@@ B-E Configuration-based Locality Optimizations
Assume that additional locality features (e.g. package locality [8]) are integrated into the scheduler. In this case, if the attacker provides the same specifications as the victim, he/she can increase the probability of having function hosts placed on the machine with a victim host, i.e.
@@ C-B Claim: The algorithm minimizes co-location of users and respects resource constraints
Preliminaries: Let $C(u)$ represent set of servers user $u$ occupies, and let $C^{*}(u)$ denote the set of hosts after assigning function $f_{u}$ to host $h$. The co-location metric is defined as:
@@
    Case 2: If no such host $h\in C(u)$ exists, the algorithm selects a host h* that satisfies the resource constraint and minimizes the increase in the co-location metric. Formally, the host is chosen as:
@@ 2. Assumption 2:
The results are shown in Figure 13. As expected, the TSC read differences for condition (1) (Same FuncApp w/ Shared Data) are significantly lower than those for condition (2) (Random FuncApps); that is, function pairs likely to be co-located indeed exhibit similar TSC readings, whereas random function pairs show much higher TSC read variances.
For accurate co-location measurement, we set the threshold to $t_{h}=10^{8}$ (indicated by the dotted line in Figure 13). Any two measurements with a difference below this threshold are considered to originate from co-located function hosts.

### v3

@@ 30 Million requests per year (Coca-Cola’s vending machine service [39]);
    Each function host is allocated 2GB of memory and 1GB of storage;
In Table II, we provide a comparison of AWS and Azure. For AWS, we provide results from the US-East-1 region. We use the AWS Lambda Pricing Calculator [10] to estimate the cost of serverless services, and their Elastic Compute Cloud (EC2) dedicated host pricing information [9] to estimate the cost of having a dedicated host. Similarly, for Azure, we use their dedicated host pricing list [12] and their pricing calculator [13] in the West US region.
@@
Using a dedicated host not only incurs substantially higher costs but can also lead to scalability limitations. For example, in AWS, a single dedicated host can accommodate only 16 a1-medium instances (based on the capacity information provided in [9]). This level of concurrency is insufficient to handle bursts of requests. In contrast, Double-Dip, as a defense integrated within the serverless framework, preserves the platform’s ability to scale elastically and can therefore support applications with high and variable demand.
@@ X-A Serverless Cloud Scheduling
Scheduler is considered one of the critical components of a serverless cloud system. In recent years, there have been works that target different features of serverless functions and propose corresponding optimizations to scheduling algorithms. Fuerst et al. [19] propose an algorithm named Consistent Hashing with Random Load Updates (CH-RLU), which improves the classic consistent hashing algorithm and takes server loads, function cold-start overheads, etc., into consideration to make trade-offs between server loads and function locality. Kaffes et al. [23] offer a taxonomy of serverless scheduling algorithms and summarize a design parameter space of these schedulers. The authors then propose their algorithm, named Hermod, which significantly improves the performance of serverless systems.
Regarding locality optimizations, Aumala et al. [8] propose PASch, which also uses consistent hashing and allows serverless systems to utilize the package locality of running functions to accelerate the function start process. Abdi et al. [1] propose to use locality hints in schedulers that specify a function’s preferences in data locality and optimize the performance of serverless functions. Li et al. [28] propose an algorithm to avoid function cold-starts in serverless systems by recycling idle function hosts of other functions and avoiding creating new function hosts.
@@ X-B Co-Location Attacks
The most relevant work is [47]. The authors propose a method to fingerprint physical servers and identify the auto-scaling feature of the Google Cloud Run platform, using which they manage to achieve instance co-location. However, compared to our work, their major focus is on the server fingerprinting step, and they only proved that their method works on one type of cloud scheduler. Our work focuses on the systemization of the co-location attack process, and our contribution is the construction of a universal attack strategy that can efficiently target different kinds of serverless cloud schedulers. Also, we propose a low-overhead mitigation scheduling algorithm, which is based on “soft” user isolation, to defend against co-location attacks.
@@ XII Ethics Considerations
The study of co-location itself does not directly compromise cloud users’ instances. It would require subsequent attacks to cause real damage, which are beyond the scope of this work. We disclosed our findings to Microsoft, although they did not classify the issue as a vulnerability, noting that side-channel defenses are already in place. These defenses, however, are orthogonal to our work, which focuses on the scheduler side.
@@ B-A Basics
The placement results of schedulers that consider only resource availability can be approximated as random events in a classical probability model. This can be explained as follows: since cluster information is agnostic to all users, from a user’s perspective, the resource availability on physical servers is a random variable following the same distribution. This symmetry indicates that the probability of a user’s instance being placed on any given node is equally $1/N$, where $N$ is the size of the cluster.
@@ B-B Random Scheduling
We first consider the scenario where every invocation to functions is randomly dispatched to nodes in the serverless cluster. In this scenario, each function invocation is considered independently, and no locality optimization is taken. If the assigned machine does not have a corresponding warm function host running, it results in a cold-start process. This scheduling policy is load-balanced but suffers from low performance due to the lack of locality optimizations.
@@
i.e., co-location tends to happen on almost every physical server if the number of invocations to victim and attacker functions is sufficiently large.
@@ B-C Invocation Locality
Now, let’s consider the scenario where the scheduler introduces optimization to avoid cold-start and always assigns function executions to existing function hosts if there are any. In this scenario, we can define $\alpha$ as the number of different attack functions and $\beta$ as the number of launched victim function hosts. In this scenario, we have $\beta\equiv 1$. As there are $\alpha$ different functions, there will be $\alpha$ different function hosts that are placed on physical machines in the cluster.
@@ B-E Configuration-based Locality Optimizations
Assume that additional locality features (e.g., package locality [8]) are integrated into the scheduler. In this case, if the attacker provides the same specifications as the victim, he/she can increase the probability of having function hosts placed on the machine with a victim host, i.e.,
@@ C-B Claim: The algorithm minimizes co-location of users and respects resource constraints
Preliminaries: Let $C(u)$ represent the set of servers user $u$ occupies, and let $C^{*}(u)$ denote the set of hosts after assigning function $f_{u}$ to host $h$. The co-location metric is defined as:
@@
    Case 2: If no such host $h\in C(u)$ exists, the algorithm selects a host $h^{*}$ that satisfies the resource constraint and minimizes the increase in the co-location metric. Formally, the host is chosen as:
@@ 2. Assumption 2:
The results are shown in Fig. 13. As expected, the TSC read differences for condition (1) (Same FuncApp w/ Shared Data) are significantly lower than those for condition (2) (Random FuncApps); that is, function pairs likely to be co-located indeed exhibit similar TSC readings, whereas random function pairs show much higher TSC read variances.
For accurate co-location measurement, we set the threshold to $t_{h}=10^{8}$ (indicated by the dotted line in Fig. 13). Any two measurements with a difference below this threshold are considered to originate from co-located function hosts.